OpenStudy (ronlover101):

PLEASE HELP (MEDAL) x − y = 5 3x + 2y = −1 (a) (2, −3) (b) (6, 1

OpenStudy (anonymous):

I'm going to assume there are more options, yes?

OpenStudy (andrewishere):

im not really understanding the question?

OpenStudy (andrewishere):

oh i see the coordinates must work for both equations

OpenStudy (ronlover101):

Determine whether each point is a solution to the system of linear equations. 1. x − y = 5 3x + 2y = −1 (a) (2, −3) (b) (6, 1)

OpenStudy (andrewishere):

it is not a or b. choice a does not work for any of them and b only works for the first equation

OpenStudy (andrewishere):

unfortunately, i can still only see choices a and b.

OpenStudy (ronlover101):

becuse those are the only choices

OpenStudy (andrewishere):

oh wait

OpenStudy (andrewishere):

choice a has a -3

OpenStudy (andrewishere):

so it works for the first equation

OpenStudy (anonymous):

\[\large \sf 6-6 \neq -1\]

OpenStudy (andrewishere):

but it equals 0 for the second equation

OpenStudy (andrewishere):

i am sorry they both work for the first equation but not for the second equation.

OpenStudy (ronlover101):

ok?

OpenStudy (anonymous):

are you supposed to solve the system \[x − y = 5 \\3x + 2y = −1\]?

OpenStudy (anonymous):

double the first one \[2x-2y=10\\ 3x+2y=-1\] then add and the \(y\) terms will go bye bye leaving \[5x=9\]

OpenStudy (ronlover101):

ok

OpenStudy (anonymous):

making \[x=\frac{9}{5}\]

OpenStudy (anonymous):

plug that in to \(x-y=5\) and solve \[\frac{9}{5}-y=5\]

OpenStudy (anonymous):

if one answer choices is not \[(\frac{9}{5},-\frac{16}{5})\] then there is a typo somewhere

OpenStudy (ronlover101):

@iambatman Please help.

OpenStudy (anonymous):

Hi, it seems you're in good hands what's the problem?

OpenStudy (ronlover101):

I can't find out out to do this problem. And I am not understanding the the other people are saying.

OpenStudy (anonymous):

Well the idea here is to get a point (x,y) so you have two equations and you need to solve for both x and y, so we have a system of equations.

OpenStudy (ronlover101):

ok

OpenStudy (anonymous):

There are a few ways to do this, substitution, elimination, matrices etc. But lets do elimination, so our goal is to first solve for either x or y. x-y=5 3x+2y=-1 So as mentioned in the name elimination, lets eliminate one of the variables. As satellite mentioned it would be a good idea to double the first equation meaning multiply it by 2, so we get \[2 \times (x-y=5) \implies 2x-2y=10\] so our system now is \[2x-2y=10\]\[3x+2y=-1\] now add these two equations what happens?

OpenStudy (ronlover101):

I don't really know what happens next.

OpenStudy (anonymous):

What is 2x+3x?

OpenStudy (ronlover101):

1?

OpenStudy (ronlover101):

I will be rightback

OpenStudy (anonymous):

2x+3x=5x right?

OpenStudy (anonymous):

witch one do you need help with?

OpenStudy (ronlover101):

YES IT DOES. @iambatman